OpenStudy (anonymous):

Integrate 6x/(sin^2x) from pi/6 , pi/2

OpenStudy (alekos):

why don't you plug in the values and see what result you get?

OpenStudy (anonymous):

how to do integral of 6x/(sin^2x)

OpenStudy (anonymous):

x is not constant how we can take it out....

hero (hero):

Use should use uv substitution for this

OpenStudy (anonymous):

ok ..thanks

OpenStudy (alekos):

sorry, didn't read your question properly. i agree with Hero, use integration by substitution

ganeshie8 (ganeshie8):

yes by parts, or if u knw complex numbers, consider complexifying....

Parth (parthkohli):

\[I:=\int \dfrac{6x}{\sin^2x}dx\]\[= \int 6x \cdot \dfrac{1}{\sin^2x}dx\]\[= \int 6x\cdot \csc^2(x)dx\] \[u = 6x \Rightarrow du = 6dx\]\[v = -\cot(x) \Rightarrow dv = \csc^2(x) dx\Rightarrow \] Now,\[I=uv - \int vdu\]\[=-6x\cot x -\int -6\cot(x)dx\]\[= -6x\cot x +6\ln|\sin(x)|+C\]

Parth (parthkohli):

I have no idea why it took so long to submit.

Parth (parthkohli):

Anyway, that's the antiderivative.